OpenStudy (anonymous):

The radius of a circle is 10/pi. How many inches long is its circumference? F. 20/pi G. 200/pi H. 10 J. 20 K. 40

OpenStudy (campbell_st):

the formula is \[C = 2 \pi r \] C = circumference r = radius just substitute and evaluate

OpenStudy (anonymous):

so it's 2*pi*10 in this case?

OpenStudy (campbell_st):

isn't it \[C = 2 \times \pi \times \frac{10}{\pi}\] which needs to be evaluted